4. A nylon rope used by mountaineers elongates 1.10 m under the weight of a 65.0-kg climber. If the rope is 45.0 m in length and 7.0 mm in diameter, what is Young's modulus for nylon?

1
Expert's answer
2022-06-06T08:55:04-0400
"\sigma=\varepsilon E"

"F/A=E\Delta l/l_0"

"E=l_0 F/(A\Delta l)\\
=45.0*65.0*9.81/(3.14*(3.5*10^{-3})^2*1.10)\\
=6.78*10^8\:\rm Pa"